How many moles of S are in each of the following quantities?

- 20 g of S

- 128 g of SO2

- 2.0 moles of Al2S3
Answer #1

1)

Molar mass of S = 32.07 g/mol


mass(S)= 20 g

use:
number of mol of S,
n = mass of S/molar mass of S
=(20 g)/(32.07 g/mol)
= 0.6236 mol
Answer: 0.62 mol

2)
Molar mass of SO2,
MM = 1*MM(S) + 2*MM(O)
= 1*32.07 + 2*16.0
= 64.07 g/mol


mass(SO2)= 128 g

use:
number of mol of SO2,
n = mass of SO2/molar mass of SO2
=(1.28*10^2 g)/(64.07 g/mol)
= 2.00 mol

1 mol of SO2 has 1 mol of S
So,
Mol of S = mol of SO2
= 2.00 mol
Answer: 2.00 mol

3)
1 mol of Al2S3 has 3 mol of S
So,
Mol of S = 3*mol of Al2S3
= 3*2.0 mol
= 6.0 mol
Answer: 6.0 mol